OpenStudy (anonymous):

Zac's weekly wage is $627.20 a) what does he earn for 4 weeks. b) Zac receives an annual leave loading of 17 1/2% on 4 weeks holiday pay. Find the amount of the annual leave loading. c) What is Zac's total holiday pay. i need working out please but simple

OpenStudy (anonymous):

a\[627.20*4=2308.8\]

OpenStudy (anonymous):

Thankyou what about b

OpenStudy (anonymous):

17 1/2% = 17.5% right

OpenStudy (anonymous):

i made a mistake the answer to a is 2508.8. sorry

OpenStudy (anonymous):

yes but what is b and c it is hard

OpenStudy (anonymous):

yes the answer is 2508. but i still dont get what is b and c please

OpenStudy (anonymous):

to find B you need to take the percentage amount and in this case (17.5%) and there are multiple ways you can turn the percentage into a decimal (move the decimal two places to the left. .175 and multiple that by the total pay) or multiply by 17.5/100 to get 439.04 to get c add those two number together 2508.8 +439.04 =2947.80